OpenStudy (haileyd):

How do i find the inverse of [-1 3] -4 2

OpenStudy (mathmale):

What methods have you learned for finding the inverse of a 2x2 matrix? There is a formula for this purpose. But you could also find the inverse using the Gauss-Jordan Method (which involves row operations). Which method is your teacher using?

OpenStudy (haileyd):

So i learned that you multiply the -1 and 4 and subtract the answer from that when you multiply 3 and -4. After that i thought would become that denominator so like 1(1/10) but that must be way off because my answer was wrong.

OpenStudy (mathmale):

That method is the "formula method" and applies only to 2 x 2 matrices. Your given matrix is a 2 x 2 matrix. What are the values of a, b, c and d?

OpenStudy (haileyd):

a=-1 b=3 c=-4 d=2 and then i would do ad-bc would be the start right?

OpenStudy (mathmale):

yes. show all work, please.

OpenStudy (haileyd):

that was what i did and got 1/10 and multiplied it by all the a b c and d values and my answer was wrong.

OpenStudy (mathmale):

Be patient, please. How did you get 1/10? I need to see your calculations.

OpenStudy (haileyd):

so -1(2)-3(-4)=10 then it becomes 1/10 and multiply it be the a,b,c,d

OpenStudy (sshayer):

\[\left| A \right|=\left|\begin{matrix}-1 & 3 \\ -4 & 2\end{matrix}\right|=(-1)2-(-4)3=-2+12=10\] co-factors of first row are 2,-(-4) i.e.,2,4 co-factors of second row are -(3),+(-1) i.e.,-3,-1 \[adjoint~ A =\left[\begin{matrix}2 & 4 \\ -3 & -1\end{matrix}\right]'=\left[\begin{matrix}2 & -3 \\ 4 & -1\end{matrix}\right]\] \[A ^{-1}=\frac{ 1 }{ 10 }\left[\begin{matrix}2 & -3 \\ 4 & -1\end{matrix}\right]\]

OpenStudy (haileyd):

i got it thank you